# DAG-aware Transformer for Causal Effect Estimation

### v1

@@ 3.3 Hyperparameter Tuning and Model Selection
Hyperparameter tuning and model selection present unique challenges in causal inference, as unlike traditional machine learning tasks with observed labels and cross-validation procedures, we cannot directly observe counterfactual potential outcomes (Saito and Yasui, 2020; Mahajan et al., 2024). To address this issue, several methods have been developed (Nie and Wager, 2017; Saito and Yasui, 2020). In our work, we adopt the approach proposed by (Saito and Yasui, 2020), which aligns well with our training scheme and offers a straightforward implementation.
@@
    We train a plug-in estimator $\hat{\tau}$ on the validation set using generalized random forests (Wager and Athey, 2018).
@@
Formally, we select the optimal estimator $\tilde{\tau}^{*}$ according to $\tilde{\tau}^{\ast}=\operatorname*{arg\,min}_{\tilde{\tau}\in\mathcal{T}}\,\text{NRMSE}\left(\hat{\tau},\tilde{\tau}\right)$, where $\mathcal{T}$ is the set of candidate estimators, and Normalized Root Mean Squared Error (NRMSE) is defined as $\text{NRMSE}=\sqrt{\frac{\frac{1}{n-1}\sum_{i=1}^{n}(\hat{\tau}(X_{i})-\tilde{\tau}(X_{i}))^{2}}{\hat{V}(\hat{\tau}(X))}}$. Here, $\{\tilde{\tau}(X_{i})\}_{i=1}^{n}$ is a set of ATE or CATE predictions by $\tilde{\tau}(\cdot)$, and $\hat{V}(\hat{\tau}(X))$ is the empirical variance of the ground-truth ATE or CATE approximated by $\hat{\tau}(\cdot)$.
@@ 4 Experiments
We evaluate our proposed method on three diverse datasets. First, we use the LaLonde datasets to assess Average Treatment Effect (ATE) estimation, and the ACIC dataset for Conditional Average Treatment Effect (CATE) estimation. Both these experiments assume causal assumptions listed in Appendix A. The specifications of baseline models for the LaLonde and ACIC experiments are provided in Appendix B.1
We also extend our evaluation to setting amid unmeasured confounding using proximal inference described in Tchetgen et al. (2024). Detailed results of the proximal inference experiment are presented in Appendix C.3.
@@ 4.1 Lalonde dataset for ATE estimation
The LaLonde datasets, derived from the National Supported Work (NSW) Demonstration program, include the Current Population Survey (CPS) and Panel Study of Income Dynamics (PSID) (LaLonde, 1986). We define the treatment (A) as participation in the NSW program, the outcome (Y) as earnings in 1978, and the covariates (X) as age, education, race, marital status, earnings in 1974, and earnings in 1975. The LaLonde CPS dataset comprises 15,992 control units and 185 treated units, while the LaLonde PSID dataset consists of 2,490 control units and the same 185 treated units. The true ATE for both datasets is 1,794.34. To ensure robust evaluation, we create 10 distinct samples from each dataset using bootstrap with replacement with different random seeds. Figures 3 and 3 present the Normalized Root Mean Squared Error (NRMSE) results with standard errors for the LaLonde CPS and PSID datasets, respectively. Tables 2 and 2 in Appendix B.2 provide the corresponding numerical values.
@@
Our proposed AIPW method achieves the lowest NRMSE on both datasets, with 48.2% and 23.2% reductions compared to AIPW (GRF) for LaLonde CPS and PSID, respectively. The performance gap is more pronounced in the LaLonde CPS dataset, suggesting better generalization to diverse control populations. Notably, our method consistently achieves lower NRMSE with smaller standard errors, indicating more accurate and stable causal effect estimations in challenging real-world scenarios. We chose not to use the RealCause simulated LaLonde data (Neal et al., 2020) due to significant discrepancies in the true ATE, which could lead to misleading comparisons.
@@ 4.2 ACIC dataset for CATE estimation
The ACIC dataset from the 2016 Atlantic Causal Inference Conference data challenge (Dorie et al., 2019) is based on real covariates with synthetically simulated treatment assignment and potential outcomes. We analyze 10 instances from different data-generating processes, each containing 58 pre-treatment variables, a binary treatment assignment, observed outcome, and ground truth potential outcomes. Figure 4 presents the NRMSE with standard errors with respect to the 10 instances of datasets for each method.
Figure 4: Mean NRMSE with Standard Error for ACIC Dataset
Our proposed AIPW method demonstrates superior performance, achieving the lowest NRMSE of 0.894, closely followed by AIPW (GRF) with 0.920. The IPW methods show higher error rates, with our IPW variant outperforming IPW (GRF) and IPW (Naive). Importantly, our proposed methods (both IPW and AIPW variants) consistently outperform their GRF counterparts, underscoring the effectiveness of our approach in CATE estimation tasks. Table 3 in Appendix B.3 provide the corresponding numerical values.
@@ 5 Conclusion
In this paper, we introduced a novel transformer-based approach for causal inference that addresses key limitations in existing methods. Our model’s primary innovation lies in its ability to encode any causal DAGs into the attention mechanism, allowing it to handle a wide range of causal scenarios. The causal-aware attention mechanism we developed explicitly models the encoded causal structure, leading to more accurate estimation of treatment effects. Our experimental results demonstrate the effectiveness of our approach across various synthetic and real-world datasets, showing improved performance compared to existing methods. While our work represents a significant step forward in causal inference using transformer architectures, there are several directions for future research. While our current approach demonstrates effectiveness for the causal graphs studied, we recognize the importance of swift adaptability to a wider range of causal structures. Future work can explore developing a more generalized encoding mechanism that can quickly accommodate diverse causal graphs without requiring extensive retraining. This could involve creating a meta-learning framework that learns to adapt to new causal structures efficiently. Additionally, we acknowledge the need to investigate the robustness of our approach against potential noise or misspecifications in the input DAG. Future studies can systematically introduce perturbations to the causal graph to assess how our model’s performance degrades under various levels of DAG uncertainty. This analysis will provide insights into the model’s resilience and help identify areas where the attention mechanism might be made more robust to structural noise.
@@ Appendix A Causal Assumptions
    This assumption implies that, conditional on the observed confounders $X$, the potential outcomes $Y^{a}$ are independent of the treatment assignment $A$. In other words, after controlling for $X$, there are no unmeasured confounders that affect both the treatment assignment and the outcome. This is also known as the "no unmeasured confounding" assumption.
@@
Appendix B Experiments assuming unconfoundedness
B.1 Baseline models
Our baseline models include Naive Inverse Probability Weighting (IPW), which uses uniform weights, and IPW with Generalized Random Forests (GRF) [Wager and Athey, 2018] for propensity score estimation. Additionally, we consider Augmented IPW (AIPW) with GRF, which employs a doubly robust estimation approach. All baseline models (except the naive model) were fine-tuned using the grf package in R.
B.2 LaLonde Dataset Results
Table 1: Results on LaLonde CPS Dataset
Method NRMSE SE of NRMSE
IPW (Naive) 5.106 0.340
IPW (GRF) 6.342 1.227
IPW (Ours) 4.209 0.414
AIPW (GRF) 1.596 0.294
AIPW (Ours) 0.826 0.194
Table 2: Results on LaLonde PSID Dataset
Method NRMSE SE of NRMSE
IPW (Naive) 9.280 0.168
IPW (GRF) 9.408 1.108
IPW (Ours) 8.857 0.118
AIPW (GRF) 2.517 0.242
AIPW (Ours) 1.934 0.177
B.3 ACIC Dataset Results
Table 3: Results on ACIC Dataset
Method NRMSE SE of NRMSE
IPW (Naive) 5.473 1.421
IPW (GRF) 2.342 0.637
IPW (Ours) 1.669 0.378
AIPW (GRF) 0.920 0.082
AIPW (Ours) 0.894 0.033
Appendix C Proximal Inference
C.1 Preliminaries
In proximal inference [Tchetgen et al., 2024], we aim to estimate the expected potential outcome $\mathbb{E}[Y^{a}]$ for each treatment level $a$, in the presence of unobserved confounders $U$, given a set of proxies $(W,Z)$ and observed confounders $X$. The key assumptions are:
Assumption 1
Assumption 2
Assumption 3
[…]
The expected potential outcomes are given by:
$\mathbb{E}[Y^{a}]=\mathbb{E}_{W,X}[h(a,W,X)]$ (7)
C.2 Methods
Building upon Kompa et al. [2022], we use our DAG-aware transformer to estimate the bridge function. Our model comprises:
  1. 1.
    A transformer encoder with DAG-masked self-attention layers, ensuring nodes attend only to their causal parents and themselves.
  2. 2.
    A Multilayer Perceptron (MLP) that combines the encoder output with raw inputs (treatment $A$ and proxies $W$) to preserve information potentially lost during encoding.
The model output $\hat{Y}$ estimates the bridge function $\hat{h}$. We compute its empirical mean for causal effect estimation, employing the state-of-the-art U-statistics kernel from Kompa et al. [2022].
C.3 Experiment
Detailed description of the Demand dataset can be found at Kompa et al. [2022]. The goal is to estimate the effect of airline ticket price $A$ on sales $Y$, confounded by unobserved demand $U$. We use fuel cost $Z$ as a treatment-inducing proxy and website views $W$ as an outcome-inducing proxy (Figure 5). We train our model on simulated datasets with sample sizes ranging from 1,000 to 10,000. Performance is evaluated using causal mean squared error (c-MSE) across 10 equally-spaced price points between 10 and 30, comparing estimated potential outcomes $\hat{E}[Y^{a}]$ against Monte Carlo simulations of the true $E[Y^{a}]$. We use a heldout dataset of 1,000 draws from $W$ to compute predictions. We perform 20 replicates for each sample size. Table C.3 summarizes the c-MSE distribution across sample sizes. Our model consistently outperforms the previous state-of-the-art model (MLP) reported in Kompa et al. [2022].
@@ C.3 Experiment
Table 4: Demand Median (IQR) values
Method Training set size
1000 5000 10000
MLP 23.68 (8.02) 16.21 (10.55) 14.25 (4.46)
Ours 10.69 (14.72) 7.67 (6.70) 5.56 (6.72)

### v2

@@ 3.3 Hyperparameter Tuning and Model Selection
Hyperparameter tuning and model selection present unique challenges in causal inference, as unlike traditional machine learning tasks with observed labels and cross-validation procedures, we cannot directly observe counterfactual potential outcomes (Saito & Yasui, 2020; Mahajan et al., 2024). To address this issue, several methods have been developed (Nie & Wager, 2017; Saito & Yasui, 2020). In our work, we adopt the approach proposed by (Saito & Yasui, 2020), which aligns well with our training scheme and offers a straightforward implementation.
@@
    We train a plug-in estimator $\hat{\tau}$ on the validation set using generalized random forests (Wager & Athey, 2018).
@@
Formally, we select the optimal estimator $\tilde{\tau}^{*}$ according to $\tilde{\tau}^{\ast}=\operatorname*{arg\,min}_{\tilde{\tau}\in\mathcal{T}}\,\text{NRMSE}\left(\hat{\tau},\tilde{\tau}\right)$, where $\mathcal{T}$ is the set of candidate estimators, and Normalized Root Mean Squared Error (NRMSE) is defined as $\text{NRMSE}=\sqrt{\frac{\frac{1}{n-1}\sum_{i=1}^{n}(\hat{\tau}(X_{i})-\tilde{\tau}(X_{i}))^{2}}{\hat{V}(\hat{\tau}(X))}}$. Here, $\{\tilde{\tau}(X_{i})\}_{i=1}^{n}$ is a set of ATE or CATE predictions by $\tilde{\tau}(\cdot)$, and $\hat{V}(\hat{\tau}(X))$ is the empirical variance of the ground-truth ATE or CATE approximated by $\hat{\tau}(\cdot)$. A detailed description of the hyper-parameter tuning process, including all considered value ranges, is available in Appendix C.
@@ 4 Experiments
We evaluate our proposed DAG-aware Transformer method on four diverse datasets, assessing its performance across different causal inference tasks and estimation methods. We use the LaLonde datasets for Average Treatment Effect (ATE) estimation. The ACIC dataset is employed for Conditional Average Treatment Effect (CATE) estimation, allowing us to assess our model’s ability to capture heterogeneous treatment effects. Additionally, we extend our evaluation to proximal inference scenarios using the Demand dataset described in (Kompa et al., 2022), testing our model’s adaptability to more complex causal structures. The causal assumptions for each of these experiments are detailed in Appendix A.
We implement and compare three main causal estimators using our DAG-aware Transformer architecture: Standardization, which estimates the outcome model directly; Inverse Probability Weighting (IPW), which focuses on estimating propensity scores; and Augmented Inverse Probability Weighting (AIPW), a doubly robust method that combines both outcome modeling and propensity score estimation. For each dataset, we compare the performance of these estimators to assess which approach works best in different scenarios. This comparison allows us to evaluate the flexibility and effectiveness of our DAG-aware Transformer across various causal inference methodologies.
For the AIPW estimator, we conduct an additional experiment to compare two training strategies: separate training, where we train the outcome model and propensity score model independently using our DAG-aware Transformer, and joint training, where we train both models simultaneously within a single DAG-aware Transformer architecture. This comparison aims to determine whether the joint learning of outcome and propensity models offers advantages over separate training in terms of estimation accuracy or computational efficiency. We hypothesize that the joint training approach may benefit from shared representations and potentially capture more nuanced interactions between confounders, treatment, and outcomes.
4.1 Evaluation metrics
We use normalized root mean square error (NRMSE) to compare our estimated ATE with the true ATE for LaLonde CPS, LaLonde PSID datasets, and the ACIC dataset (described below). This metric provides a standardized measure of the estimation accuracy, allowing for comparisons across different datasets and methods. In the case of the Demand dataset and proximal inference, we follow the evaluation protocol outlined in (Kompa et al., 2022). Specifically, performance is evaluated using causal mean squared error (c-MSE) across 10 equally-spaced price points between 10 and 30, comparing estimated potential outcomes $\hat{E}[Y^{a}]$ against Monte Carlo simulations of the true $E[Y^{a}]$. This approach allows us to assess the model’s ability to estimate counterfactual outcomes across a range of treatment values.
4.2 Baseline models
To evaluate the performance of our DAG-aware Transformer, we compare it against several established baseline models, each chosen for their relevance to causal inference tasks. Our primary baselines include Generalized Random Forests (GRF) (Wager & Athey, 2018), a versatile non-parametric method that adapts random forests for heterogeneous treatment effect estimation, and Multilayer Perceptron (MLP), a standard neural network architecture that serves as a comparison to our more complex transformer-based model. Additionally, for the Inverse Probability Weighting (IPW) estimator, we include a Naive IPW baseline that uses uniform weights, demonstrating the importance of proper weighting in causal inference.
For the Lalonde and ACIC datasets, we implement GRF and MLP for each estimator (G-formula, IPW, and AIPW), with the naive IPW additionally used for the IPW estimator. In the case of the Demand dataset, we employ MLP as our primary baseline due to its effectiveness in handling the dataset’s specific characteristics (Kompa et al., 2022). To ensure a fair comparison, all baseline models (except the naive IPW) underwent careful hyperparameter tuning.
@@ 4.3 Lalonde dataset for ATE estimation
The LaLonde datasets, derived from the National Supported Work (NSW) Demonstration program, include the Current Population Survey (CPS) and Panel Study of Income Dynamics (PSID) (LaLonde, 1986). We define the treatment (A) as participation in the NSW program, the outcome (Y) as earnings in 1978, and the covariates (X) as age, education, race, marital status, earnings in 1974, and earnings in 1975. The LaLonde CPS dataset comprises 15,992 control units and 185 treated units, while the LaLonde PSID dataset consists of 2,490 control units and the same 185 treated units. The true ATE for both datasets is 1,794.34. To ensure robust evaluation, we create 10 distinct samples from each dataset using bootstrap with replacement with different random seeds. Figures 2 and 3 present the Normalized Root Mean Squared Error (NRMSE) results with standard errors for the LaLonde CPS and PSID datasets, respectively. Table 1 in Appendix B.1 provide the corresponding numerical values.
@@
Our DAG-aware Transformer model consistently outperformed the baseline methods across both datasets and all estimation techniques. For the Lalonde-CPS dataset, in G-formula estimation, our Transformer achieved the lowest NRMSE of 0.901, compared to 0.925 for GRF and 0.922 for MLP. The Transformer’s performance in IPW was equally impressive, with an NRMSE of 0.614, significantly outperforming GRF (6.342) and slightly better than MLP (0.631). In AIPW, our jointly trained Transformer-Joint model achieved the best performance with an NRMSE of 0.351, substantially outperforming all other methods including the separately trained Transformer-Sep (1.512).
Similar trends were observed in the Lalonde-PSID dataset. For G-formula, our Transformer model achieved the lowest NRMSE of 0.964, compared to 1.009 for GRF and 0.979 for MLP. In IPW estimation, the Transformer (NRMSE 2.198) outperformed both GRF (9.408) and MLP (2.258). For AIPW, the Transformer-Joint model again showed superior performance with an NRMSE of 0.922, significantly better than GRF (2.517), MLP (1.249), and Transformer-Sep (3.237).
Notably, the jointly trained Transformer model (Transformer-Joint) in AIPW consistently outperformed its separately trained counterpart (Transformer-Sep) across both datasets, highlighting the benefits of joint training in capturing complex interactions between confounders, treatment, and outcomes. The superior performance of joint training in ATE estimation can be attributed to several factors. Firstly, joint training allows for shared representation learning, potentially capturing complex interactions between confounders, treatment, and outcomes that might be missed when models are trained separately (Shalit et al., 2017). This shared learning can be particularly beneficial when the same features influence both treatment assignment and outcomes. Secondly, joint training can lead to more efficient use of the data, especially in smaller datasets, by leveraging information across tasks (Künzel et al., 2019). Furthermore, the ATE, being a population-level estimate, may benefit from the regularizing effect of joint training, which can prevent overfitting to noise in individual components (Shi et al., 2019). Joint training also allows for end-to-end optimization of the entire causal inference pipeline, potentially leading to a better global optimum (Yao et al., 2018). Additionally, in scenarios where there’s model misspecification, joint training might offer robustness by allowing complementary strengths of the outcome and propensity models to compensate for each other’s weaknesses (Chernozhukov et al., 2018). These advantages seem to outweigh the potential benefits of specialized modeling in the ATE context, where the goal is to estimate an average effect rather than heterogeneous effects across subpopulations.
@@ 4.4 ACIC dataset for CATE estimation
The ACIC dataset from the 2016 Atlantic Causal Inference Conference data challenge (Dorie et al., 2019) is based on real covariates with synthetically simulated treatment assignment and potential outcomes. We analyze 10 instances from different data-generating processes, each containing 58 pre-treatment variables, a binary treatment assignment, observed outcome, and ground truth potential outcomes. Figure 4 presents the NRMSE with standard errors with respect to the 10 instances of datasets for each method. Table 2 in Appendix B.2 provide the corresponding numerical values.
Figure 4: Mean NRMSE with Standard Error for LaLonde PSID Dataset
Our DAG-aware Transformer model demonstrated competitive performance in CATE estimation across all estimation techniques on the ACIC dataset. In the G-formula estimation, our Transformer achieved the lowest NRMSE of 0.344, slightly outperforming GRF (0.346) and showing a substantial improvement over MLP (0.417). This suggests that our model effectively captures the underlying causal structure in estimating conditional treatment effects.
For the IPW estimator, all methods showed higher NRMSE values, indicating the challenging nature of this approach for CATE estimation. Nevertheless, our Transformer model achieved the best performance with an NRMSE of 4.102, marginally better than MLP (4.142) and GRF (4.281). This consistent, albeit small, improvement across different estimation techniques highlights the robustness of our approach.
In the AIPW setting, both variants of our Transformer model outperformed the baseline methods, with the separately trained Transformer (Transformer-Sep) achieving the lowest NRMSE of 0.698, slightly outperforming the jointly trained version (Transformer-Joint) with an NRMSE of 0.814. This contrasts with our ATE estimation results, where joint training was superior. This phenomenon aligns with recent literature on doubly robust estimators for CATE. (Nie & Wager, 2017) argue that jointly fitting the outcome and propensity models can lead to overfitting in CATE estimation, particularly when the propensity model is misspecified. They propose separate estimation of nuisance functions to maintain doubly robust properties. Similarly, (Kennedy, 2020) suggests that separate estimation can help avoid bias amplification in CATE settings. The complexity of capturing heterogeneous treatment effects in CATE may benefit from the specialization afforded by separate training, allowing each model to focus on its specific task without interference. These findings underscore the importance of considering both joint and separate training approaches in causal inference tasks, as the optimal strategy may depend on the specific estimation problem and dataset characteristics.
4.5 Demand for proximal inference
The goal of this experiment is to estimate the effect of airline ticket price $A$ on sales $Y$, confounded by unobserved demand $U$. We use fuel cost $Z$ as a treatment-inducing proxy and website views $W$ as an outcome-inducing proxy (Figure 6 in Appendix B.3). We train our model on simulated datasets with sample sizes ranging from 1,000 to 50,000. We use a heldout dataset of 1,000 draws from $W$ to compute predictions. We perform 20 replicates for each sample size. Figure 5 summarizes the c-MSE distribution across sample sizes. Table  3 in Appendix B.3 provides the corresponding numerical values.
Figure 5: Median (IQR) of c-MSE for Demand Dataset
Our Transformer model consistently outperformed the MLP baseline, with the most significant improvements observed when using the NMMR-U estimator. For instance, with a training set size of 10,000, our NMMR-U Transformer achieved a median c-MSE of 5.56 (IQR: 6.72), substantially lower than the MLP’s 14.25 (IQR: 4.46). This performance advantage was maintained across all training set sizes, from 1,000 to 50,000 samples.
@@ 5 Conclusion
In this paper, we introduced a novel transformer-based approach for causal inference that addresses key limitations in existing methods. Our model’s primary innovation lies in its ability to encode any causal DAGs into the attention mechanism, allowing it to handle a wide range of causal scenarios. The causal-aware attention mechanism we developed explicitly models the encoded causal structure, leading to more accurate estimation of treatment effects. Our experimental results demonstrate the effectiveness of our approach across various synthetic and real-world datasets, showing improved performance compared to existing methods. While our work represents a step forward in causal inference using transformer architectures, there are several directions for future research. First, we recognize the importance of swift adaptability to a wider range of causal structures. Future work can explore developing a more generalized encoding mechanism that can quickly accommodate diverse causal graphs without requiring extensive retraining. This could involve creating a meta-learning framework that learns to adapt to new causal structures efficiently. Additionally, we acknowledge the need to investigate the robustness of our approach against potential noise or misspecifications in the input DAG. Future studies can systematically introduce perturbations to the causal graph to assess how our model’s performance degrades under various levels of DAG uncertainty.
@@ Appendix A Causal Assumptions
    This assumption implies that, conditional on the observed confounders $X$, the potential outcomes $Y^{a}$ are independent of the treatment assignment $A$. In other words, after controlling for $X$, there are no unmeasured confounders that affect both the treatment assignment and the outcome. This is also known as the ”no unmeasured confounding” assumption.
@@
Appendix B Experiments
B.1 LaLonde Dataset Results
Table 1: NRMSE for ATE estimation on LaLonde datasets
Dataset Estimator Method Mean NRMSE SE NRMSE
Lalonde-CPS G-formula GRF 0.925 0.002
MLP 0.922 0.006
Transformer 0.901 0.005
IPW GRF 6.342 1.227
MLP 0.631 0.104
Transformer 0.614 0.100
AIPW GRF 1.596 0.294
MLP 0.517 0.142
Transformer-Sep 1.512 0.578
Transformer-Joint 0.351 0.130
Lalonde-PSID G-formula GRF 1.009 0.021
MLP 0.979 0.004
Transformer 0.964 0.014
IPW GRF 9.408 1.108
MLP 2.258 0.745
Transformer 2.198 0.705
AIPW GRF 2.517 0.242
MLP 1.249 0.259
Transformer-Sep 3.237 0.571
Transformer-Joint 0.922 0.264
B.2 ACIC Dataset Results
Table 2: NRMSE for CATE estimation on ACIC dataset
Estimator Method Mean NRMSE SE NRMSE
G-formula GRF 0.346 0.044
MLP 0.417 0.062
Transformer 0.344 0.054
IPW GRF 4.281 1.388
MLP 4.142 1.278
Transformer 4.102 1.272
AIPW GRF 0.857 0.059
MLP 1.304 0.112
Transformer-Sep 0.698 0.057
Transformer-Joint 0.814 0.052
B.3 Proximal inference
@@
Table 3: Median c-MSE (IQR) for Demand dataset across different training set sizes
Method Model 1000 5000 10000 50000
NMMR-V MLP 23.41 (11.26) 30.74 (17.73) 42.88 (29.45) 62.18 (16.97)
Transformer 21.54 (17.42) 24.46 (17.93) 21.37 (10.12) 27.50 (16.30)
NMMR-U MLP 23.68 (8.02) 16.21 (10.55) 14.25 (4.46) 14.27 (12.47)
Transformer 10.69 (14.72) 7.67 (6.70) 5.56 (6.72) 6.51 (5.90)
Appendix C Hyperparameter Tuning
We performed extensive hyperparameter tuning for each dataset and estimation method. The hyperparameter spaces explored are detailed below. For each configuration, we report the range of values considered during the tuning process.
Table 4: Hyperparameter tuning ranges for Lalonde-CPS dataset
Parameter G-formula IPW AIPW
Number of epochs 80 20 20
Batch size 32 32 32
Learning rate 0.001 0.001 0.001
L2 penalty 3e-05, 3e-03 3e-05, 3e-03 3e-05, 3e-03
Network width (MLP) 80 80 80
Input layer depth (MLP) 2–4 1–2 2–6
Number of layers (encoder) 2–4 1–2 1–2
Dropout rate 0.0001 0.0001–0.001 0.0001
Embedding dimension (encoder) 40 40 40
Feedforward dimension (encoder) 80 80 80
Number of heads (encoder) 2 1–2 1–2
Encoder weight (alpha) 0.02 0.002–0.02 0.02
Table 5: Hyperparameter tuning ranges for Lalonde-PSID dataset
Parameter G-formula IPW AIPW
Number of epochs 100 30 30
Batch size 32, 64 64 32, 64
Learning rate 0.001–0.01 0.001 0.001
L2 penalty 3e-05 3e-05 3e-05–3e-03
Network width (MLP) 80 10, 20 40
Input layer depth (MLP) 6–16 1 4–8
Number of layers (encoder) 1–2 1 1–2
Dropout rate 0.0001 0.0001 0.0001
Embedding dimension (encoder) 40 10–40 20–40
Feedforward dimension (encoder) 80 20–80 40–80
Number of heads (encoder) 1–2 1 2
Encoder weight (alpha) 0.002–0.02 0.02–0.2 0.02–0.2
Table 6: Hyperparameter tuning ranges for ACIC dataset
Parameter G-formula IPW AIPW
Number of epochs 500 30 500
Batch size 64 64, 128, 256 64, 128, 256
Learning rate 1e-03 1e-03 1e-04
L2 penalty 3e-08 3e-05 3e-08, 3e-04
Network width (MLP) 40 40, 60, 80 40, 60, 80, 120
Input layer depth (MLP) 8–16 4–16 2–16
Number of layers (encoder) 16 2–4 2–8
Dropout rate 0.0001 0.0001 0.0001–0.003
Embedding dimension (encoder) 256 40, 60, 80 256
Feedforward dimension (encoder) 1024 80, 320 512, 1024
Number of heads (encoder) 4 2 2–4
Encoder weight (alpha) 0.02–0.2 0.002–0.2 0.1–2
Table 7: Hyperparameter tuning ranges for Demand dataset (Proximal Inference)
Parameter Range
Number of epochs 1000
Batch size 32, 64
Learning rate 0.001
L2 penalty 3e-06
Network width (MLP) 160
Input layer depth (MLP) 8, 16
Number of layers (encoder) 1, 2
Dropout rate 0
Embedding dimension (encoder) 40
Feedforward dimension (encoder) 40, 80
Number of heads (encoder) 1, 2
Encoder weight (alpha) 0.001–0.025
For each dataset and estimation method, we performed a grid search over these hyperparameter spaces. The final model for each configuration was selected based on the best performance on a held-out validation set.